# Improved composable key rates for CV-QKD

### v3

@@ I Introduction
The paper is structured as follows. In Sec. II we derive our general formula for the secret key rate of a generic CV-QKD protocol; this is done by refining previous theory and adopting a number of improvements, including a different approach to tensor-product reduction after error correction (proven in Appendix A). In Sec. III, we apply the results to relevant examples of CV-QKD protocols, showing the improvements in terms of key rate with respect to previous literature. Sec. IV is for conclusions.
@@ II.1 Output state of a CV-QKD protocol
Consider a CV-QKD protocol where $N$ single-mode systems are transmitted from Alice $A$ to Bob $B$. A portion $n$ of these systems will be used for key generation, while a portion $m=N-n$ will be used for parameter estimation. Let us assume that the bosonic communication channel depends on a number $n_{\text{pm}}$ of parameters $\mathbf{p}=(p_{1},p_{2},\ldots)$ (e.g., transmissivity and thermal noise of the channel). These parameters are estimated by the parties and we will account for their partial knowledge at the end of the derivations. For now, let us assume that Alice and Bob has perfect knowledge of $\mathbf{p}$.
@@
[…]
@@ II.2 Error correction and epsilon correctness
Then, as a verification, Alice and Bob publicly compare hashes computed over $\mathbf{l}$ and $\hat{\mathbf{l}}$. If these hashes coincide, the two parties go ahead with the probability $p_{\text{ec}}$, otherwise, they abort the protocol. We denote by $T_{\text{ec}}$ the case of a successful verification (no abort), so that $\rho_{|T_{\text{ec}}}$ represents a conditional post-EC state. More specifically, the hash comparison requires Bob to send $\left\lceil-\log_{2}\varepsilon_{\text{cor}}\right\rceil$ bits to Alice for some suitable $\varepsilon_{\text{cor}}$ (the number of these bits is typically small in comparison to $\mathrm{leak}_{\text{ec}}$). Parameter $\varepsilon_{\text{cor}}$ is called the $\varepsilon$-correctness [32, Sec. 4.3] and it bounds the probability that Alice’s and Bob’s corrected sequences are different even if their hashes coincide. The probability of such an error is bounded by [33]
$p_{\text{ec}}\mathrm{Prob}(\hat{\mathbf{l}}\neq\mathbf{l}|T_{\text{ec}})\leq p_{\text{ec}}2^{-\left\lceil-\log_{2}\varepsilon_{\text{cor}}\right\rceil}\leq\varepsilon_{\text{cor}}.$ (6)
@@ II.3 Equivalence to a projection process
$\Gamma=\{(\mathbf{k},\mathbf{l}):\mathrm{Prob}(\hat{\mathbf{l}}\neq\mathbf{l}|T_{\text{ec}})\leq\varepsilon_{\text{cor}}\},$ (8)
@@
$\cal{E}_{\mathrm{guess}}(\left|\mathbf{k},\mathbf{l}\right\rangle\left\langle\mathbf{k},\mathbf{l}\right|)=|\hat{\mathbf{l}},\mathbf{l}\rangle\langle\hat{\mathbf{l}},\mathbf{l}|,$ (10)
@@ II.6 Leftover hash bound
we certainly realize the secrecy bound in Eq. (16). If we also impose the condition of correctness in Eq. (6), we reach the condition of epsilon security for Alice and Bob expressed by Eq. (17). Setting $\varepsilon_{\text{h}}:=\varepsilon_{\text{sec}}-\varepsilon_{\text{s}}$ and re-arranging Eq. (25), we derive the following upper-bound for the binary klength of the key (converse leftover hash bound)
@@ II.10 Achievable key rate for optimal PA
$s_{n}^{\text{opt}}\geq H_{\text{min}}^{\varepsilon_{\text{s}}}(B^{n}|E^{n})_{\sigma^{n}}+2\log_{2}(\sqrt{2}\varepsilon_{\text{h}}),$ (42)
We can see that the $-1$ difference between Eqs. (26) and (42) become an extra $-p_{\text{ec}}/N$ in Eq. (41). Because $N$ is typically large, we also see that $R_{\text{LB}}\simeq R_{\text{UB}}$.
@@ II.11.2 Formula for Gaussian-modulated coherent state protocols
[…]
@@ III.2 Gaussian modulation of coherent states with heterodyne detection
Figure 2: Improved composable secret key rate [upper bound of Eq. (40)] for the Gaussian modulated coherent-state protocol with homodyne detection (blue solid line) and heterodyne detection (red solid line) with respect to the block size $N$. These lines coincide with those computed from the lower bound of Eq. (41). The corresponding dashed lines are computed using Eq. (43), based on previous literature. Loss is set to $7$ dB, while all the other parameters are chosen as in Fig. 1.
@@ III.3 Parameter estimation and final performance
$\sigma_{\xi}$ $=\sqrt{\frac{2}{V_{0}m}}\frac{\eta T\xi+V_{0}+u_{\text{del}}}{\eta T_{\text{wc}}}.$ (88)
In the equations above, $V_{0}=1$ is for homodyne detection and $V_{0}=2$ is for heterodyne detection. Then, in Eq. (87), the term $c_{\mathrm{PE}}$ can be set to zero [14] (in fact, another choice would be $c_{\mathrm{PE}}=2$ [15] based on a weaker assumption [39]). The parameter $w$ that connects the worst-case values with $\varepsilon_{pe}$ is simply given by an inverse error function when we assume a Gaussian approximation for the parameters, i.e.,
@@
Thus, by using $\hat{\mathbb{p}}=(\hat{T},\hat{\xi})$ and $\mathbb{p}_{\text{wc}}=(T_{\text{wc}},\xi_{\text{wc}})$, we compute the PE rate as in Eq. (51) to be replaced in Eq. (40) for both the homodyne and heterodyne protocols. Let us assume ad hoc values for $p_{\text{ec}}$ and $\beta$ (the exact numerical values of these parameters are known after a realistic implementation or simulation of EC, as discussed in Sec. II.15). Then, we show the performances of the two protocols in Figs. 1 and 2. More specifically, in Fig. 1, we depict the secret key rate versus channel loss, while, in Fig. 2, we show its behavior with respect to block size. For the sake of comparison, we have also included the results based on previous literature [20, 21] (refined in Sec. II.10). From the figures, we can see a significant improvement in the key rate performance both in terms of robustness to loss and smaller block size.
IV Conclusions
@@
defined over a reduced alphabet $\Gamma\subseteq\mathcal{X^{\prime}}\otimes\mathcal{X}$ for the classical system $C^{\prime}C$ and a subsequent guess channel $\cal{E}_{\mathrm{guess}}$ applied to $C^{\prime}C$ [cf. Eqs. (9) and (10) in the main text]. Then due to the monotonicity of the purified distance under completely positive trace non-increasing maps, i.e., projections, CPTP maps, and partial trace operations [37, Theorem 3.4], we have
@@ B.1 Estimating the transmissivity
$\text{Var}(\widehat{T})$ $=\frac{4V_{\text{Cov}}C^{2}_{x,y}}{\eta^{2}(\sigma_{x}^{2})^{4}}=\frac{4C^{4}_{x,y}}{\eta^{2}(\sigma_{x}^{2})^{4}}\frac{\left[c_{\text{pe}}+\frac{\sigma_{z}^{2}}{\eta T\sigma_{x}^{2}}\right]}{V_{0}m}$ (131)
@@
as in Eq. (87) of the main text up to replacing $V=\sigma^{2}_{z}$. For $m\gg 1$, we may assume that the distribution of $\widehat{T}$ is Gaussian with variance given by Eq. (131). Therefore, we may write that
@@
$P[\widehat{p}\geq p+w\sigma_{T}]\leq\varepsilon_{\text{pe}}.$ (137)
@@
$P\left[\widehat{p}-p\geq+w\sigma_{p}\right]\leq\varepsilon_{\text{pe}}$
$P\left[\frac{\widehat{p}-p}{\sigma_{p}}\geq w\right]\leq\varepsilon_{\text{pe}}.$ (138)
@@
[…]
@@ B.2 Estimating the noise
The sum above follows a central chi-squared distribution with $d_{\text{f}}=V_{0}m$ and, therefore, with mean $V_{0}m$ and variance $2V_{0}m$. Then $\widehat{\sigma}_{z}^{2}$ follows the same distribution but rescaled by $\frac{\sigma_{z}^{2}}{V_{0}m}$. Thus, its mean value is $\sigma_{z}^{2}$ while its variance $V_{z}=\frac{2(\sigma_{z}^{2})^{2}}{V_{0}m}$. From this, we may write
$[\sigma_{z}^{2}]_{\text{wc}}\simeq\sigma_{z}^{2}+w\sqrt{V_{z}}$ (152)
@@
$=$ $\sigma_{z}^{2}+\sigma_{z}^{2}\frac{\sqrt{2}}{\sqrt{V_{0}m}}\sqrt{2\ln\varepsilon_{\text{pe}}}+\mathcal{O}\left(\frac{1}{m}\right)$
$\simeq$ $\sigma_{z}^{2}+\sigma_{z}^{2}\frac{\sqrt{2}}{\sqrt{V_{0}m}}\sqrt{2\ln\varepsilon_{\text{pe}}},$ (153)

### v5

@@ I Introduction
The paper is structured as follows. In Sec. II we derive our general formula for the secret key rate of a generic CV-QKD protocol; this is done by refining previous theory and adopting a number of improvements, including a different approach to tensor-product reduction after error correction (proven in Appendix A). In Sec. III, we apply the results to relevant examples of CV-QKD protocols, showing the improvements in terms of key rate with respect to previous literature. Sec. IV contains some clarifications and Sec. V is for conclusions.
@@ II.1 Output state of a CV-QKD protocol
Consider a CV-QKD protocol where $N$ single-mode systems are transmitted from Alice $A$ to Bob $B$. A portion $n$ of these systems will be used for key generation, while a portion m=Nnm=N-n will be used for parameter estimation. Let us assume that the bosonic communication channel depends on a number $n_{\text{pm}}$ of parameters $\mathbf{p}=(p_{1},p_{2},\ldots)$ (e.g., transmissivity and thermal noise of the channel). These parameters are estimated by the parties and we will account for their partial knowledge at the end of the derivations. For now, let us assume that Alice and Bob has perfect knowledge of $\mathbf{p}$.
@@
In the case of a protocol based on the Gaussian modulation of coherent states, the multi-ary symbols are the output of analog-to-digital conversion (ADC) from Alice’s and Bob’s quadratures, $x$ and $y$, i.e., we have $x\overset{\text{ADC}}{\rightarrow}k$ and $y\overset{\text{ADC}}{\rightarrow}l$. If the protocol is based on the homodyne detection, we have that $y$ is randomly created by a random switching between the $q$ and $p$ quadrature (with Alice choosing the corresponding quadrature for each instance, upon Bob’s classical communication). If the protocol is based on the heterodyne detection, both $q$ and $p$ quadratures are used, so we have $y=(q_{B},p_{B})\overset{\text{ADC}}{\rightarrow}(l_{q},l_{p})$ followed by the concatenation $l=l_{q}l_{p}$ so that the dimension is $d=d_{q}+d_{p}$, where $d_{q}$ ($d_{p}$) refers to the dimension of $l_{q}$ ($l_{p}$). Finally, in the case of CV-QKD protocols based on discrete-alphabet coherent states, no ADC is necessary and the discretized variables are directly expressed by the encoding variables.
@@ II.2 Error correction and epsilon correctness
Then, as a verification, Alice and Bob publicly compare hashes computed over $\mathbf{l}$ and $\hat{\mathbf{l}}$. If these hashes coincide, the two parties go ahead with the probability $p_{\text{ec}}$, otherwise, they abort the protocol. We denote by $T_{\text{ec}}$ the case of a successful verification (no abort), so that $\rho_{|T_{\text{ec}}}$ represents a conditional post-EC state. More specifically, the hash comparison requires Bob to send $\left\lceil-\log_{2}\varepsilon_{\text{cor}}\right\rceil$ bits to Alice for some suitable $\varepsilon_{\text{cor}}$ (the number of these bits is typically small in comparison to $\mathrm{leak}_{\text{ec}}$). Parameter $\varepsilon_{\text{cor}}$ is called the $\varepsilon$-correctness [32, Sec. 4.3] and it bounds the probability that Alice’s and Bob’s corrected sequences are different. The probability of such an error is bounded by [33]
$p_{\text{ec}}\mathrm{Prob}(\hat{\mathbf{l}}\neq\mathbf{l}|T_{\text{ec}})\leq\varepsilon_{\text{cor}}.$ (6)
@@ II.3 Equivalence to a projection process
$\Gamma=\{(\mathbf{k},\mathbf{l}):\mathrm{Prob}(\mathbf{k}\neq\mathbf{l})\leq\varepsilon_{\text{cor}}\},$ (8)
@@
[…]
@@ II.6 Leftover hash bound
we certainly realize the secrecy bound in Eq. (16). If we also impose the condition of correctness in Eq. (6), we reach the condition of epsilon security for Alice and Bob expressed by Eq. (17). Setting $\varepsilon_{\text{h}}:=\varepsilon_{\text{sec}}-\varepsilon_{\text{s}}$ and re-arranging Eq. (25), we derive the following upper-bound for the binary length of the key (converse leftover hash bound)
@@ II.10 Achievable key rate for optimal PA
$s_{n}^{\text{opt}}\geq H_{\text{min}}^{\varepsilon_{\text{s}}}(B^{n}|E^{n})_{\sigma^{n}}+2\log_{2}(\sqrt{2}\varepsilon_{\text{h}}).$ (42)
We can see that the 1-1 difference between Eqs. (26) and (42) become an extra $p_{\text{ec}}/N$ in Eq. (41). Because $N$ is typically large, we also see that $R_{\text{LB}}\simeq R_{\text{UB}}$.
@@ II.11.2 Formula for Gaussian-modulated coherent state protocols
where $I(x:y)\geq I(k:l)$ is Alice and Bob’s mutual information computed over their continuous variables. Second, we exploit the data processing inequality for Eve’s Holevo bound, so $\chi(l:E)_{\rho}\leq\chi(y:E)_{\rho}$ under digitalization $y\overset{\text{ADC}}{\rightarrow}l$. Thus, we can use the asymptotic rate
@@ III.2 Gaussian modulation of coherent states with heterodyne detection
Figure 2: Improved composable secret key rate [upper bound of Eq. (40)] for the Gaussian modulated coherent-state protocol with homodyne detection (blue solid line) and heterodyne detection (red solid line) with respect to the block size $N$. These lines coincide with those computed from the lower bound of Eq. (41). The corresponding dashed lines are computed using Eq. (43), based on previous literature. Loss is set to 77 dB, while all the other parameters are chosen as in Fig. 1.
@@ III.3 Parameter estimation and final performance
$\sigma_{\xi}$ $=\sqrt{\frac{2}{V_{0}m}}\frac{\eta T\xi+V_{0}+u_{\text{el}}}{\eta T_{\text{wc}}}.$ (88)
In the equations above, $V_{0}=1$ is for homodyne detection and $V_{0}=2$ is for heterodyne detection. Then, in Eq. (87), the term $c_{\mathrm{pe}}$ can be set to zero [14] (in fact, another choice would be $c_{\mathrm{pe}}=2$ [15] based on a weaker assumption [39]). The parameter $w$ that connects the worst-case values with $\varepsilon_{\mathrm{pe}}$ is simply given by an inverse error function when we assume a Gaussian approximation for the parameters, i.e.,
@@
[…]
Figure 3: Rate and epsilon security, from block to session.
IV Discussion
Let us clarify some important points about the epsilon security and the secret key rate. As previously discussed, under stable channel conditions, one can consider a QKD session of $n_{\mathrm{bks}}$ blocks. Each block of size $N$ undergoes the three subsequent processes depicted in Fig. 3. These are the following:
  • PE, affected by errors $\varepsilon_{\mathrm{pe}}$ and $\varepsilon_{\mathrm{ent}}$;
  • EC, performed with success probability $p_{\mathrm{ec}}$, and affected by error $\varepsilon_{\mathrm{cor}}$;
  • PA, affected by error $\varepsilon_{\mathrm{sec}}$.
The PE errors $\varepsilon_{\mathrm{pe}}$ and $\varepsilon_{\mathrm{ent}}$ are introduced before EC, so they will appear in the final epsilon security $\varepsilon$ with probability $p_{\mathrm{ec}}$ (with probability $1-p_{\mathrm{ec}}$ the block is not processed and therefore there is no forward-propagated PE error). Thus, on average, the two types of PE errors contribute as $p_{\mathrm{ec}}\varepsilon_{\mathrm{pe}}$ and $p_{\mathrm{ent}}\varepsilon_{\mathrm{ent}}$ in the final $\varepsilon$.
Situation is different for the correctness $\varepsilon_{\mathrm{cor}}$. This is related to the joint probability that the verification step of EC succeeds and the strings are different [cf. Eq. (6)]. This means that it is already defined as an average (unconditional) quantity. In fact, we may write that
$\mathrm{Prob}(T_{\text{ec}}\mathrm{~and~}\hat{\mathbf{l}}\neq\mathbf{l})=p_{\text{ec}}\mathrm{Prob}(\hat{\mathbf{l}}\neq\mathbf{l}|T_{\text{ec}})\leq\varepsilon_{\text{cor}}$ (91)
is equivalent to the unconditional probability
$\mathrm{Prob}(\hat{\mathbf{l}}\neq\mathbf{l})=p_{\text{ec}}\mathrm{Prob}(\hat{\mathbf{l}}\neq\mathbf{l}|T_{\text{ec}})$
$+(1-p_{\text{ec}})\mathrm{Prob}(\hat{\mathbf{l}}\neq\mathbf{l}|\mathrm{not~}T_{\text{ec}})\leq\varepsilon_{\text{cor}}.$ (92)
Similarly, the secrecy $\varepsilon_{\mathrm{sec}}$ is the joint probability that the protocol succeeds (i.e., passes EC) and the output key is not secure [cf. Eq. (16)]. Therefore, this is also an average (unconditional) quantity that contributes directly as is to the epsilon security $\varepsilon$.
For this reason, the epsilon security takes the form of Eq. (71), i.e.,
$\varepsilon=\varepsilon_{\text{cor}}+\varepsilon_{\text{sec}}+p_{\text{ec}}\varepsilon_{\text{ent}}+p_{\text{ec}}n_{\text{pm}}\varepsilon_{\text{pe}}.$ (93)
However, this expression assumes the knowledge of $p_{\mathrm{ec}}$, so it is preferable to use the simpler bound
$\varepsilon\leq\varepsilon_{\text{cor}}+\varepsilon_{\text{sec}}+\varepsilon_{\text{ent}}+n_{\text{pm}}\varepsilon_{\text{pe}}.$ (94)
For the session, the total epsilon security is $\varepsilon_{\mathrm{ses}}=n_{\mathrm{nks}}\varepsilon$.
For the secret key rate, we may consider the conditional or unconditional secret key rate. The conditional rate $R_{\mathrm{cond}}$ is defined on a successfully-corrected block, while the unconditional rate $R$ is the average output rate (the one considered in the previous sections of the paper). Note that we may write
$R_{\mathrm{cond}}=\frac{s_{n}}{N},~R=p_{\text{ec}}R_{\mathrm{cond}},~R_{\mathrm{ses}}=R.$ (95)
Finally, another important clarification regards PE. Our theory does not depend on how estimators ($\hat{T}$ and $\hat{\xi}$) and worst-case values ($T_{\mathrm{wc}}$ and $\xi_{\mathrm{wc}}$) are calculated. In fact, for any QKD session, these quantities can be sampled and computed directly from the available experimental data. For theoretical investigations, one can adopt the derivations presented in Appendix B but other estimation methods can be employed.
V Conclusions
@@
Acknowledgements
This work was supported by the EPSRC via the UK Quantum Communications Hub (Grant No. EP/T001011/1).
@@ Acknowledgements
[…]
@@ Appendix B Details on parameter estimation for Gaussian-modulated protocols
Note that, in this appendix, we adopt a different notation for the heterodyne protocol. In the main text, $y$ represented both Bob’s quadratures, i.e., $y=(q_{B},p_{B})$. Here, $y$ represents Bob’s generic quadrature, i.e. $y=q_{B}$ or $p_{B}$.
@@ B.1 Estimating the transmissivity
For the central limit theorem (CLT), the quantity $\frac{\widehat{C}_{x,y}}{\sqrt{V_{\text{Cov}}}}$ tends to a Gaussian. In fact, we can bound the Kolmogorov distance of its actual distribution from a Gaussian distribution. For typically-large block sizes, one can use the Berry-Esseen inequality to check, numerically, that the distance becomes quickly negligible.
@@
[…]
@@
as in Eq. (87) of the main text up to replacing $V=\sigma^{2}_{x}$.
For large enough $m\gg 1$, we have again a good convergence in the CLT and, therefore, we may assume that the distribution of $\widehat{T}$ becomes Gaussian with variance given by Eq. (136). Therefore, we may write that
@@
$\mathrm{Prob}[\widehat{p}\geq p+w\sigma_{T}]\leq\varepsilon_{\text{pe}}.$ (142)
@@
$\mathrm{Prob}\left[\widehat{p}-p\geq w\sigma_{p}\right]\leq\varepsilon_{\text{pe}}$
$\mathrm{Prob}\left[\frac{\widehat{p}-p}{\sigma_{p}}\geq w\right]\leq\varepsilon_{\text{pe}}.$ (143)
@@
of the normal variable $\frac{\widehat{p}-p}{\sigma_{p}}$. We use its connection to the error function $\text{erf}(.)$ to write
@@ B.2 Estimating the noise
The sum above follows a central chi-squared distribution with $d_{f}=V_{0}m$ and, therefore, with mean $V_{0}m$ and variance $2V_{0}m$. Then $\widehat{\sigma}_{z}^{2}$ follows the same distribution but rescaled by $\frac{\sigma_{z}^{2}}{V_{0}m}$. Thus, its mean value is $\sigma_{z}^{2}$ while its variance $V_{z}=\frac{2(\sigma_{z}^{2})^{2}}{V_{0}m}$. From this, we may write
$\sigma_{z}^{2}]_{\text{wc}}\simeq\sigma_{z}^{2}+w\sqrt{V_{z}}$ (157)
@@
$=$ $\sigma_{z}^{2}+\sigma_{z}^{2}\frac{\sqrt{2}}{\sqrt{V_{0}m}}\sqrt{2\ln\varepsilon_{\text{pe}}^{-1}}+\mathcal{O}\left(\frac{1}{m}\right)$
$\simeq$ $\sigma_{z}^{2}+\sigma_{z}^{2}\frac{\sqrt{2}}{\sqrt{V_{0}m}}\sqrt{2\ln\varepsilon_{\text{pe}}^{-1}},$ (158)